OpenStudy (anonymous):

how many real roots does the function f(x)=x^2-3x+4 have? A. –7 B. 0 C. 1 D. 2

OpenStudy (anonymous):

lol surely not-7

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

solve \[x^2-3x+4=0\] and see

OpenStudy (anonymous):

or since you are not actually asked for the solutions, compute \[b^2-4ac\] with \(a=1,b=-3,c=4\)

OpenStudy (anonymous):

once you see that the answer is negative, that means there are no solutions

OpenStudy (anonymous):

yeah i equaled it out to be one half something

OpenStudy (anonymous):

obviously 2 real roots

OpenStudy (anonymous):

so that means its 0??

OpenStudy (anonymous):

i believe its 0

OpenStudy (anonymous):

yeah \[(-3)^2-4\times 1\times 4=9-16=\text{some negative number}\]

OpenStudy (anonymous):

so no real solutions for sure

OpenStudy (anonymous):

thank you @satellite73

OpenStudy (anonymous):

yw

OpenStudy (triciaal):